OpenStudy (anonymous):

find the sum 3/4 + 1/6

OpenStudy (akshay_budhkar):

now you try this 1

OpenStudy (akshay_budhkar):

i will check if its right

OpenStudy (anonymous):

3/4 + 1/6= 9/12 + 2/12 =11/12